OpenStudy (anonymous):

Suppose you divide a polynomial by a binomial. How do you know if the binomial is a factor of the polynomial? Create a example problem that has a binomial which IS a factor of the polynomial being divided, and another problem that has a binomial which is NOT a factor of the polynomial being divided.

OpenStudy (anonymous):

The binomial is a factor if the remainder is zero after division.

OpenStudy (anonymous):

Remainder theorem: \[\frac{p(x)}{f(x)}=q(x)+r(x)\] i.e. a polynomial divided by another function equals a quotient function plus a remainder function. If the remainder equals zero, then f(x) is a factor of p(x). \[\rightarrow p(x)=q(x)f(x)\]

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

pick whatever functions you want for q(x), f(x), and r(x), then work backwards to get the polynomial dividend, p(x).

OpenStudy (anonymous):

Here's an example from arithmetic: \[108\div9=12\] 9 is a factor of 108, so there is no remainder. Here is the same statement using algebra: \[(x^2+x-2) \div (x-1)=(x+2)\] Let x=10 in the above, and you'll see it's exactly the same. :-)